### v2

# Large Deviation Principles for Abelian Monoids

Abstract.
Following work of Mehrdad and Zhu [MZ16] and of Liu [Liu04a, Liu04b], we prove a large deviation principle for a broad class of integer-valued additive functions defined over abelian monoids. As a corollary, we obtain a large deviation principle for a generalized form of the Erdős-Kac theorem due to Liu.
1. Introduction
The celebrated Erdős-Kac Theorem [EK40] says that if $\omega(n)$ is the number of distinct prime factors of a positive integer $n$, then
$\frac{\omega(n)-\log\log n}{\sqrt{\log\log n}}$ (1)
is normally distributed with mean 00 and standard deviation 1 (cf. Theorem 1.2).
Much work has been done on the rate at which (1) converges to $N(0,1)$, see [RT58] for one such example. Likewise, one can study the tails of the distribution as the upper bound for $n$ grows. Indeed, in [MZ16], Mehrdad and Zhu prove a large deviation principle for a large class of “strongly additive” functions $g:\mathbb{N}\to\mathbb{N}$, including when $g$ is the number-of-distinct-prime-divisors function, $\omega$, thus giving a large deviation principle for the distributions of (1) over $n\leq X$ as $X\to\infty$.
In a different direction, Liu proves a generalization of the Erdős-Kac theorem for classes of abelian monoids $\mathcal{M}$ generated by a set $\mathcal{P}$ [Liu04a]. This general framework includes, for example, the cases where $\mathcal{P}$ is the set of rational primes, the set of irreducible monic polynomials over a finite field, the set of effective 0-cycles of a nice variety over a finite field, and the case where $\mathcal{P}$ is a Beurling system of generalized primes. See [Beu37] or [DZ16] for background on the latter.
We will work in this same general setting, which we now describe. Let $\mathcal{P}$ be a set together with a “norm” map ${\mathbf{N}:\mathcal{P}\to\mathbb{N}\setminus\{1\}.}$ Let $\mathcal{M}$ be the multiplicative monoid generated by the elements of $\mathcal{P}$. That is, every element $m\in\mathcal{M}$ can be expressed as $m=\prod_{p\in\mathcal{P}}p^{e_{p}}$ where almost all of the $e_{p}\in\mathbb{Z}_{\geq 0}$ are zero. Extend $\mathbf{N}$ to a monoid morphism $\mathcal{M}\to\mathbb{N}$ by taking, for any $m\in\mathcal{M}$, ${\mathbf{N}(m)=\prod_{p\in\mathcal{P}}\mathbf{N}(p)^{e_{P}}.}$
Throughout, we make the following two assumptions on the growth rate of $\mathcal{M}$ and $\mathcal{P}$ with respect to $\mathbf{N}$.
Condition 1.1.
For $\mathcal{P},\mathcal{M},\mathbf{N}$ as above, assume
  1. (1)
    ${\displaystyle\sum_{\begin{subarray}{c}m\in\mathcal{M}\\ \mathbf{N}(m)\leq X\end{subarray}}1=aX+O(X^{b})}$ for some $a>0$ and $b\in[0,1)$, and
  2. (2)
    ${\displaystyle\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ \mathbf{N}(p)\leq X\end{subarray}}1=O\left(\frac{X}{\log X}\right)}$.
In this setting, Liu proves the following generalization of the Erdős-Kac Theorem.
Theorem 1.2 (Theorem 1, [Liu04a]).
If $\mathcal{M},\mathcal{P},\mathbf{N}$ are as above and satisfy Condition 1.1, then
[…]
Our goal is to prove a generalization of the large deviation principle in [MZ16] to the setting of abelian monoids considered by Liu. In particular, the large deviation principle in our main theorem, Theorem 1.5 below, applies to the distributions (2).
Our large deviation result will apply to a class of “strongly additive” functions.
Condition 1.3.
Let $g:\mathcal{M}\to\mathbb{N}$ be a strongly additive function, i.e. $g$ satisfies the following two conditions:
$g(p^{k})$ $=g(p)\text{ for all prime elements }p\in\mathcal{P},$
$g(mn)$ $=g(m)+g(n)\text{ for all }m,n\in\mathcal{M}\text{ such that }(m,n)=1.$
We suppose further that there exists a probability measure $\rho$ on $\mathbb{R}$ satisfying the following two conditions.
  1. (1)
    For any $\theta\in\mathbb{R}$, we have $\int_{\mathbb{R}}e^{\theta y}\rho(dy)<\infty$.
  2. (2)
    Let $A\subset\mathbb{R}$ be any Borel measurable set. Define probability measures $\rho_{X}$ given by
    $\rho_{X}(A):=\frac{\displaystyle\sum_{\begin{subarray}{c}g(p)\in A\\ \mathbf{N}(p)\leq X\end{subarray}}\frac{1}{\mathbf{N}(p)}}{\displaystyle\sum_{\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}}.$ (3)
    Then for any $\theta\in\mathbb{R}$, we have $\int_{\mathbb{R}}e^{\theta y}\rho_{X}(dy)\to\int_{\mathbb{R}}e^{\theta y}\rho(dy)$.
The application we have in mind is when $g$ is the number-of-distinct-prime-divisors function, $\omega$. On $\mathcal{M}$, this means for $m=\prod_{p\in\mathcal{P}}p^{e_{p}}$, let $\omega(m)$ be the number of $p\in\mathcal{P}$ for which $e_{p}\geq 1$.
Definition 1.4.
For each $n\in\mathbb{N}$, we denote by $V(n)$ a uniformly chosen monoid element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq n\}$. For each $p\in\mathcal{P}$, let $Z$ be a random variable such that $Z_{p}=1$ if $V(n)$ is divisible by $p$, and $Z_{p}=0$ otherwise.
We are now ready to state the main theorem.
Theorem 1.5.
Let $\mathcal{M}$ be a multiplicative abelian monoid generated by a set $\mathcal{P}$ and endowed with a norm function $\mathbf{N}:\mathcal{M}\to\mathbb{N}$, all satisfying Condition 1.1. Let $g:\mathcal{M}\to\mathbb{N}$ be an arithmetic function satisfying Condition 1.3. Let $W$ be a random variable over $\mathbb{N}$ defined as
$W(n):=g(V(n)),$ (4)
where $V(n)$ is a uniformly chosen monoidal element from the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq n\}$, see Definition 1.4. Then for any Borel measurable set $A\subset\mathbb{R}$, the probability $\mathbb{P}\left[\frac{W(X)}{\log\log X}\in A\right]$ satisfies a large deviation principle with speed $\log\log X$ and rate function
$I(x):=\sup_{\theta\in\mathbb{R}}\left\{\theta x-\int_{\mathbb{R}}(e^{\theta y}-1)\rho(dy)\right\}.$
More explicitly, we have for any Borel measurable set $A\subset\mathbb{R}$,
$-\inf_{x\in A^{o}}I(x)$ $\leq\liminf_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\frac{W(X)}{\log\log X}\in A\right]$
$\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\frac{W(X)}{\log\log X}\in A\right]\leq-\inf_{x\in\overline{A}}I(x),$
where $A^{o}$ is the interior of $A$, and $\overline{A}$ is the closure of $A$.
The strategy of the proof will closely follow that employed in [MZ16], but with suitable modifications that accommodate the more general setting.
2. Applications
Note that taking $\mathcal{P}$ to be the set of rational primes and $\mathbf{N}(p)=p$, Theorem 1.5 recovers exactly Theorem 2 of Mehrdad and Zhu [MZ16]. Furthermore, taking $\mathcal{P}$ to be the set of monic, irreducible polynomials $f$ over a finite field $\mathbb{F}_{q}$ and $\mathbf{N}(f)=q^{\deg f}$ recovers the large deviation result of Feng, Wang, and Yang [FWY20].
Two additional settings are natural to consider for applying Theorem 1.5 in conjunction with Theorem 1.2:
  • Number Fields: Let $\mathcal{O}_{K}$ be the ring of integers of a number field, $K$. Let $\mathcal{P}$ be the set of prime ideals of $\mathcal{O}_{K}$, so $\mathcal{M}$ is the set of ideals of $\mathcal{O}_{K}$. As usual, for $\mathfrak{p}\in\mathcal{P}$, let $\mathbf{N}(\mathfrak{p})=|\mathcal{O}_{K}/\mathfrak{p}|$. Condition 1.3 is satisfied by work of Weber [Web96], see also [Lan94, Chapter XIII] (for I), and by the prime ideal theorem (for II).
  • Nice Varieties: Let $V$ be any $d$-dimensional smooth, projective irreducible, geometrically integral variety defined over a finite field, $\mathbb{F}_{q}$. Let $\mathcal{P}$ be the set of closed points of $V/\mathbb{F}_{q}$, so $\mathcal{M}$ is the monoid of effective 00-cycles. For $P\in\mathcal{P}$, take $\mathbf{N}(P)=q^{d\cdot\deg(P)}$, where $\deg(P)$ is the size of the Galois orbit of $P$ in $V(\mathbb{\bar{F}}_{q})$. Condition 1.3 holds for this choice of $\mathcal{P},\mathcal{M},\mathbf{N}$ by [LW54]; see [Ros02] for an exposition of the case of curves over $\mathbb{F}_{q}$.
In the examples above, taking $g$ to be the appropriate number-of-distinct-prime-divisors function, $\omega$, on $\mathcal{M}$, Theorem 1.5 gives a large deviation principle with an explicit rate function for the corresponding version of the Erdős-Kac Theorem given by Theorem 1.2. We can use [MZ16, Corollary 3] to explicitly compute the rate function $I(x)$ as
$I(x):=\begin{cases}x\log x-x+1&\text{ if }x\geq 0,\\ +\infty&\text{ otherwise}.\end{cases}$ (5)
3. Preliminary Results
In this section we gather some results and prove some lemmas in preparation for the proof of Theorem 1.5.
The following version of Mertens’ second theorem will be a useful estimate.
Lemma 3.1 (Lemma 2, [Liu04b]).
If $\mathcal{P},\mathcal{M},\mathbf{N}$ satisfy Condition 1.1, then
$\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ \mathbf{N}(p)<X\end{subarray}}\frac{1}{\mathbf{N}(p)}=\log\log X+O\left(\frac{1}{\log X}\right).$
Notice Lemma 3.1 implies that the number-of-distinct-prime-divisors function $\omega$ on $\mathcal{M}$ satisfies Condition 1.3.
We next state a version of Gärtner-Ellis Theorem as given in [MZ16]; it is this result from which we will obtain the desired large deviation principle.
Theorem 3.2 (Page 152 of [MZ16], and Theorem 2.3.6 of [DZ98]).
Let $Z_{n}$ be a sequence of random variables on $\mathbb{R}$. Let $a_{n}$ be a sequence of positive numbers such that $\lim_{n\to\infty}a_{n}=\infty$. Suppose that for any $\theta\in\mathbb{R}$, the limit
$\Lambda(\theta):=\lim_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{E}[\exp\left(\theta a_{n}Z_{n}\right)]$
exists and is differentiable for every $\theta\in\mathbb{R}$. Then for any Borel measurable set $A\subset\mathbb{R}$, the probability $\mathbb{P}[Z_{n}\in A]$ satisfies a large deviation principle with speed $a_{n}$ and rate function
$I(x):=\sup_{\theta\in\mathbb{R}}\left\{\theta x-\Lambda(\theta)\right\}.$
More concretely, we have for any Borel measurable set $A\subset\mathbb{R}$,
$-\inf_{x\in A^{o}}I(x)\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]\leq\liminf_{n\to\infty}\frac{1}{a_{n}}\log\mathbb{P}[Z_{n}\in A]\leq-\inf_{x\in\overline{A}}I(x)$
For every element $p\in\mathcal{P}$, define independent random variables $Y_{p}$ for which
$Y_{p}=\begin{cases}1\quad&\text{with probability }\frac{1}{\mathbf{N}(p)}\\ 0\quad&\text{with probability }1-\frac{1}{\mathbf{N}(p)}\end{cases}.$
Lemma 3.3.
Given $p\in\mathcal{P}$, we recall the random variable $Z_{p}$ supported over the set $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$ from Definition 1.4. For distinct $p_{1},...,p_{k}\in\mathcal{P}$ we have
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]=\frac{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq\lfloor\frac{X}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor\}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}},$ (6)
and
$\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}]=\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}.$ (7)
Proof.
Given $m\in\mathcal{M}$, we have $Z_{p_{1}}(m)Z_{p_{2}}(m)\cdots Z_{p_{k}}(m)=1$ if and only if there exists an element $a\in\mathcal{M}$ such that $m=p_{1}p_{2}\cdots p_{k}a$. For such an $m$, we have $\mathbf{N}(m)\leq X$ if and only if $\mathbf{N}(a)\leq\lfloor\frac{X}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor$. Hence, we have
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\sum_{\begin{subarray}{c}m\in\mathcal{M}\\ \mathbf{N}(m)\leq X\end{subarray}}\mathbbm{1}_{p_{1}p_{2}\cdots p_{k}\mid m}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}}$
=#$=\frac{\#\{m\in\mathcal{M}:m=p_{1}p_{2}\cdots p_{k}a,\mathbf{N}(m)\leq X\}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}}$
=#$=\frac{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq\lfloor\frac{X}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor\}}{\#\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}}.$
The second equation follows immediately from the definition of $Y_{p}$. ∎
Lemma 3.4.
Suppose Condition 1.1.
  1. (1)
    For distinct primes $p_{1},\cdots,p_{k}\in\mathcal{P}$, we have
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]=\frac{1}{aX}\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
  2. (2)
    For distinct primes $p_{1},\cdots,p_{k}\in\mathcal{P}$, we have
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]\leq\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}]+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
  3. (3)
    For sufficiently large $X$ there exists an explicit positive constant $M$, independent of the choice of distinct primes $p_{1},\cdots,p_{k}\in\mathcal{P}$, such that
    $\mathbb{E}[Z_{p_{1}}Z_{p_{2}}...Z_{p_{k}}]\leq M\cdot\mathbb{E}[Y_{p_{1}}Y_{p_{2}}...Y_{p_{k}}].$
  4. (4)
    For sufficiently large $X$, there exists an explicit constant $M$ such that for any non-negative sequence of real numbers $\{\theta_{p}\}_{p\in\mathcal{P}}$ we have
    $\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Z_{p}\right)\right]\leq M\cdot\mathbb{E}\left[\exp\left(\sum_{\mathbf{N}(p)\leq X}\theta_{p}Y_{p}\right)\right].$
The implied constants of all the error terms are independent of the choice of $p_{1},\cdots,p_{k}$.
Proof.
By Condition 1.1 and Lemma 3.3, we have
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]$ $=\frac{\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\rfloor+O\left(\frac{X^{b}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right)}{aX+O(X^{b})}$
$=\frac{1}{aX}\left\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\right\rfloor+O\left(\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}\right).$
The second statement follows from using the inequality
$\frac{1}{aX}\left\lfloor\frac{aX}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}\right\rfloor\leq\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}=\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
To prove the third statement, we divide into two cases. Suppose $p_{1},p_{2},\cdots,p_{k}$ are distinct primes such that $\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})>X$. Then because all the random variables $Z_{p_{1}},Z_{p_{2}},\cdots,Z_{p_{k}}$ are supported over the probability space $\{m\in\mathcal{M}:\mathbf{N}(m)\leq X\}$, we have
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]=0<\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}=\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
Now suppose $p_{1},p_{2},\cdots,p_{k}$ are distinct primes such that $\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})\leq X$. Then we have
$\frac{X^{b-1}}{\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}$ $=\frac{1}{X^{1-b}\mathbf{N}(p_{1})^{b}\mathbf{N}(p_{2})^{b}\cdots\mathbf{N}(p_{k})^{b}}$
$\leq\frac{1}{\mathbf{N}(p_{1})\mathbf{N}(p_{2})\cdots\mathbf{N}(p_{k})}=\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
Hence, by the second statement of our lemma, there exists an absolute constant $M_{1}>0$ such that
$\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]\leq M_{1}\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}].$
We take $M:=\max(1,M_{1})$ to obtain the third statement of the lemma.
The fourth statement follows by using Taylor expansion. We crucially use the fact that for any non-negative integers $r_{1},r_{2}\cdots,r_{k}$ we have
$\mathbb{E}[Z_{p_{1}}^{r_{1}}Z_{p_{2}}^{r_{2}}\cdots Z_{p_{k}}^{r_{k}}]=\mathbb{E}[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}]\leq M\cdot\mathbb{E}[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}]=M\cdot\mathbb{E}[Y_{p_{1}}^{r_{1}}Y_{p_{2}}^{r_{2}}\cdots Y_{p_{k}}^{r_{k}}].$
The next three lemmas generalize Lemmas 7, 8, and 9 of [MZ16] for the abelian monoidal setting.
Lemma 3.5.
For any $\varepsilon>0$,
$\limsup_{C\to\infty}\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\Bigg{(}\bigg{|}\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ g(p)>C\\ \mathbf{N}p<X\end{subarray}}g(p)Z_{p}\bigg{|}\geq\varepsilon\log\log X\Bigg{)}=-\infty.$
Proof.
If
$\limsup_{C\to\infty}\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\Bigg{(}\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ g(p)>C\\ \mathbf{N}p<X\end{subarray}}g(p)Z_{p}\geq\varepsilon\log\log X\Bigg{)}=-\infty,$ (8)
and
$\limsup_{C\to\infty}\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\Bigg{(}\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ g(p)>C\\ \mathbf{N}p<X\end{subarray}}g(p)Z_{p}\leq-\varepsilon\log\log X\Bigg{)}=-\infty,$ (9)
then the conclusion of the lemma follows. We begin by proving (8).
By the exponential [^1] For an integrable random variable $X$ with finite, non-zero variance and $\varepsilon>0$, $P(X\geq\varepsilon)\leq e^{-\theta\varepsilon}\mathbb{E}(e^{\theta X})$ for any $\theta>0.$ Chebyshev’s Inequality with any $\theta>0$, and the third statement of Lemma 3.4, there exists an explicit constant $M>0$ such that
$\begin{split}&\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\geq\varepsilon\log\log X\right)\\ &\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{E}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Z_{p}\right)\right]-\theta\varepsilon\\ &\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\left(\log\mathbb{E}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]+\log M\right)-\theta\varepsilon\\ &=\limsup_{X\to\infty}\log\mathbb{E}\frac{1}{\log\log X}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]-\theta\varepsilon.\end{split}$ (10)
Applying Lemma 3.1, $\sum_{\mathbf{N}p\leq X}\frac{1}{\mathbf{N}(p)}\approx\log\log X$, and then using that $\log(x+1)\leq x$, we have
$\begin{split}&\hskip 14.0pt\limsup_{X\to\infty}\log\mathbb{E}\frac{1}{\log\log X}\left[\exp\left(\sum_{\begin{subarray}{c}g(p)>C\\ \mathrm{N(p)<X}\end{subarray}}g(p)Y_{p}\right)\right]-\theta\varepsilon\\ &\leq\limsup_{X\to\infty}\frac{\sum_{\begin{subarray}{c}g(p)>C\\ \mathbf{N}(p)\leq X\end{subarray}}\log(e^{\theta g(p)}-1)\mathbf{N}(p)^{-1}+1)}{\sum_{\mathbf{N}(p)\leq X}\mathbf{N}(p)^{-1}}-\theta\varepsilon\\ &\leq\limsup_{X\to\infty}\frac{\sum_{\begin{subarray}{c}g(p)>C\\ \mathbf{N}(p)\leq X\end{subarray}}(e^{\theta g(p)}-1)\mathbf{N}(p)^{-1}}{\sum_{\mathbf{N}(p)\leq X}\mathbf{N}(p)^{-1}}-\theta\varepsilon.\end{split}$ (11)
Finally, we apply Condition 1.3 (2) to the last term to obtain
$\limsup_{X\to\infty}\frac{\sum_{\begin{subarray}{c}g(p)>C\\ \mathbf{N}(p)\leq X\end{subarray}}(e^{\theta g(p)}-1)\mathbf{N}(p)^{-1}}{\sum_{\mathbf{N}(p)\leq X}\mathbf{N}(p)^{-1}}\leq\limsup_{X\to\infty}\int_{y>C}(e^{\theta y}-1)\rho_{X}(dy)\leq\int_{y>C}e^{\theta y-1}\rho(dy).$ (12)
Notice $\int_{y>C}e^{\theta y-1}\rho(dy)\to 0$ as $C\to\infty$. Taken together, (10), (11), and (12) show
$\limsup_{C\to\infty}\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\Bigg{(}\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ g(p)>C\\ \mathbf{N}p<X\end{subarray}}g(p)Z_{p}\geq\varepsilon\log\log X\Bigg{)}=-\theta\varepsilon.$
Since this holds for all $\theta>0$, (8) follows. The proof of (9) is nearly identical. ∎
Now, define $k_{X}:=X^{\frac{1}{(\log\log X)^{2}}}$.
Lemma 3.6.
Let $A(X,C):=\{p\in\mathcal{P}:k_{X}\leq\mathbf{N}(p)\leq X,|g(p)|\leq C\}$. For any $\epsilon>0$, we have
$\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\geq\epsilon\log\log X\right]=-\infty.$ (13)
Proof.
By Condition 1.1 and the third statement of Lemma 3.4, for any $\theta>0$ and sufficiently large $X$ there exists an explicit constant $M>0$ such that
$\log\mathbb{E}\left[\exp\left(\theta\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\right)\right]$ $\leq\log\mathbb{E}\left[\exp\left(\theta C\sum_{p\in A(X,C)}Y_{p}\right)\right]+\log M$
$=\sum_{p\in A(X,C)}\log\left((\exp\left(\theta C\right)-1)\cdot\frac{1}{\mathbf{N}(p)}+1\right)+\log M$
$\leq(\exp(\theta C)-1)\cdot\sum_{k_{X}\leq\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}+\log M.$
By Lemma 3.1, we have
$\sum_{k_{X}\leq\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}=2\log\log\log X+O\left(\frac{(\log\log X)^{2}}{\log X}\right).$
We then use exponential Chebyshev’s inequality to conclude
$\hskip 14.0pt\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{P}\left[\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\geq\epsilon\log\log X\right]$
$\leq\limsup_{X\to\infty}\frac{1}{\log\log X}\log\mathbb{E}\left[\exp\left(\theta\left|\sum_{p\in A(X,C)}g(p)Z_{p}\right|\right)\right]-\theta\epsilon$
$\leq-\theta\epsilon.$
Let $\theta$ grow arbitrarily large to conclude the lemma. ∎
Lemma 3.7.
Denote by $B(X,C):=\{p\in\mathcal{P}:\mathbf{N}(p)\leq k_{X},|g(p)|\leq C\}$. Then for any $\theta\in\mathbb{R}$, we have
$\lim_{X\to\infty}\left|\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Z_{p}\right)\right]-\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Y_{p}\right)\right]\right|=0.$ (14)
Proof.
Let $K$ be any constant. Consider the following three expressions:
$\sum_{r\leq K\log\log X}\frac{|\theta|^{r}}{r!}\left|\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Z_{p}\right)^{r}\right]-\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Y_{p}\right)^{r}\right]\right|,$ (15)
$\sum_{r>K\log\log X}\frac{|\theta|^{r}}{r!}\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Z_{p}\right)^{r}\right],$ (16)
$\sum_{r>K\log\log X}\frac{|\theta|^{r}}{r!}\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Y_{p}\right)^{r}\right].$ (17)
Then by Taylor expansion we have
[…]
By the second statement of Lemma 3.4, for any tuples of non-negative integers $r_{1},\cdots,r_{k}$ and sufficiently large $X$ there exists an absolute constant $B>0$ such that
$\left|\mathbb{E}\left[Z_{p_{1}}^{r_{1}}Z_{p_{2}}^{r_{2}}\cdots Z_{p_{k}}^{r_{k}}\right]-\mathbb{E}\left[Y_{p_{1}}^{r_{1}}Y_{p_{2}}^{r_{2}}\cdots Y_{p_{k}}^{r_{k}}\right]\right|=\left|\mathbb{E}\left[Z_{p_{1}}Z_{p_{2}}\cdots Z_{p_{k}}\right]-\mathbb{E}\left[Y_{p_{1}}Y_{p_{2}}\cdots Y_{p_{k}}\right]\right|\leq BX^{b-1}.$
Therefore, we have
$\hskip 14.0pt\left|\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Z_{p}\right)^{r}\right]-\mathbb{E}\left[\left(\sum_{p\in B(X,C)}g(p)Y_{p}\right)^{r}\right]\right|$
$\leq\left|\sum_{k=1}^{r}\sum_{r_{i}}\frac{r!}{r_{1}!r_{2}!\cdots r_{k}!}\frac{1}{k!}\sum_{p_{j}}g(p_{1})^{r_{1}}\cdots g(p_{k})^{r_{k}}\cdot\left(\mathbb{E}\left[Z_{p_{1}}^{r_{1}}Z_{p_{2}}^{r_{2}}\cdots Z_{p_{k}}^{r_{k}}\right]-\mathbb{E}\left[Y_{p_{1}}^{r_{1}}Y_{p_{2}}^{r_{2}}\cdots Y_{p_{k}}^{r_{k}}\right]\right)\right|$
$\leq\sum_{k=1}^{r}\sum_{r_{i}}\frac{r!}{r_{1}!r_{2}!\cdots r_{k}!}\frac{1}{k!}\sum_{p_{j}}\frac{B\cdot C^{r_{1}+\cdots+r_{k}}}{X^{1-b}}\leq\frac{B}{X^{1-b}}\left(\sum_{p\in B(X,C)}C\right)^{r}\leq\frac{B\cdot(C\cdot k_{X})^{r}}{X^{1-b}},$
where the $r_{i}$’s range over tuples of integers $(r_{1},\cdots,r_{k})$ such that $r_{1}+\cdots+r_{k}=r$, and the $p_{j}$’s ranges over prime elements $p_{1},\cdots,p_{k}\in B(X,C)$. By Stirling’s approximation for $r!$, there exists an absolute constant $B^{\prime}>0$ such that
[…]
$\leq B^{\prime}\cdot\exp\left(-(1-b)\log X\right)\cdot K\log\log X\cdot\left[\max_{r\leq K\log\log X}\left(\frac{e|\theta|Ck_{X}}{r}\right)^{r}\right]$
$\leq B^{\prime}\cdot\exp\left(-(1-b)\log X\right)\cdot K\log\log X\cdot\left(\frac{e|\theta|Ck_{X}}{K\log\log X}\right)^{K\log\log X}$
$\leq B^{\prime}K\cdot\exp\left(-(1-b)\log X+\log\log\log X+K(\log|\theta|+\log C)\log\log X+\frac{\log X}{\log\log X}\right).$
Hence, there exists an explicit constant $B^{\prime\prime}>0$ such that for sufficiently large $X$,
(15)$\leq B^{\prime\prime}e^{-\frac{1}{2}(1-b)\log X}.$
Note that the third statement of Lemma 3.4 implies that for sufficiently large $X$ there exists an explicit constant $M>0$ such that
[…]
so it suffices to consider (17) for sufficiently large $X$. We have
[…]
$\leq\sum_{r>K\log\log X}\frac{|\theta|^{r}}{r!}\cdot r\cdot C^{r}\cdot\left(\sum_{\mathbf{N}p\leq k_{X}}\frac{1}{\mathbf{N}p}\right)^{r}.$
By Lemma 3.1 and Stirling’s approximation for $r!$, we have
[…]
$\leq\sum_{r>K\log\log X}\exp\left(r\log|\theta|+r\log C+r+\log r-r\log r+r\log\log\log k_{X}\right).$
Because $r>K\log\log X$ inside the summation, we have
$\log r>\log K+\log\log\log X>\log K+\log\log\log k_{X}.$
Hence, there exist explicit constants $B_{*},B_{**}>0$ such that for sufficiently large $K$, we have
[…]
$\leq\sum_{r>K\log\log X}B_{*}\exp\left(-\frac{1}{2}r\log K\right)\leq B_{**}\exp\left(-\frac{1}{2}\log K\log\log X\right).$
Combining all the upper bounds for (15), (16), and (17), and choosing $K>1$, we have
00 0 $\leq\lim_{X\to\infty}\left|\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Z_{p}\right)\right]-\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Y_{p}\right)\right]\right|$
$\leq\lim_{X\to\infty}B^{\prime\prime}\cdot e^{-\frac{1}{2}(1-b)\log X}+(M+1)\cdot B_{**}\cdot e^{-\frac{1}{2}\log K\log\log X}=0.$
Remark 3.8.
Using Lemma 3.7, one can in fact show that
$\limsup_{X\to\infty}\frac{1}{\log\log X}\log\left|\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Z_{p}\right)\right]-\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Y_{p}\right)\right]\right|=-\infty.$
4. Proof
We are now ready to prove Theorem 1.5. Take
$B(X,C)=\{p\in\mathcal{P}:|g(p)|<C,\mathbf{N}(p)\leq k_{X}\}.$
For any real number $\theta$,
$\begin{split}\log\mathbb{E}\left[\exp\left(\theta\sum_{p\in B(X,C)}g(p)Y_{p}\right)\right]&=\log\prod_{p\in B(X,C)}\mathbb{E}\left[\exp(\theta g(p)Y_{P})\right]\\ &=\sum_{p\in B(X,C)}\log\left(\frac{1}{\mathbf{N}(p)}e^{\theta g(p)}+1-\frac{1}{\mathbf{N}(p)}\right).\end{split}$ (18)
For $p$ with sufficiently large $\mathbf{N}(p)$, the Taylor expansion for $\log(1+x)$ shows
$\log\left(\frac{1}{\mathbf{N}(p)}e^{\theta g(p)}+1-\frac{1}{\mathbf{N}(p)}\right)=\frac{1}{\mathbf{N}(p)}e^{\theta g(p)}-\frac{1}{\mathbf{N}(p)}+O_{C,\theta}\left(\frac{1}{\mathbf{N}(p)^{2}}\right).$ (19)
By Lemma 3.1,
$\frac{1}{\log\log X}\sum_{\mathbf{N}(p)\leq X}\frac{1}{\mathbf{N}(p)}=1\quad\text{as}\quad X\to\infty.$ (20)
Further,
$\lim_{X\to\infty}\frac{\log\log k_{X}}{\log\log X}=1.$ (21)
Using the definition of $\rho_{X}$, Condition 1.3(2) and equations (18), (19), (20), and (21) show
$\lim_{X\to\infty}\frac{1}{\log\log X}\mathbb{E}\left[\exp(\theta\sum_{p\in B(X,c)}g(p)Y_{p})\right]=\int_{-C}^{C}(e^{\theta y}-1)\rho(dy),$ (22)
if $\rho\{C\}=0$ and $\rho\{-C\}$. This latter condition is not an obstacle, since the measure $\rho$ on $\mathbb{R}$ has only countably many atoms. So we can choose infinite, increasing sequences of $C$ for which both $\rho\{C\}=0$ and $\rho\{-C\}=0$.
Lemma 3.7 and (22) imply
$\lim_{X\to\infty}\frac{1}{\log\log X}\mathbb{E}\left[\exp(\theta\sum_{p\in B(X,c)}g(p)Z_{p})\right]=\int_{-C}^{C}(e^{\theta y}-1)\rho(dy).$
By the Gärtner-Ellis Theorem (Theorem 3.2),
$\mathbb{P}\left(\frac{\sum_{p\in B(X,c)}g(p)Z_{p}}{\log\log X}\in A\right)$
satisfies a large deviation principle with rate function
$I_{C}(x)=\sup_{\theta\in\mathbb{R}}\left\{\theta x-\int_{-C}^{C}(e^{\theta y}-1)\rho(dy)\right\}.$
Taking $C\to\infty$, and using Lemmas 3.5 and 3.6 to handle the tail error estimates on $I_{C}(X)$, we have that
$\mathbb{P}\left(\frac{W(X)}{\log\log X}\in A\right)$
satisfies a large deviation principle with rate function
$I(x)=\lim_{C\to\infty}I_{C}(x)=\sup_{\theta\in\mathbb{R}}\left\{\theta x-\int_{-\infty}^{\infty}(e^{\theta y}-1)\rho(dy)\right\},$
as desired.
5. Acknowledgments
This work was completed while the first author visited the second at the Max Planck Institute for Mathematics in Bonn; we thank MPIM for its generous support and for providing excellent working conditions. We also thank Sachin Kumar for pointing out a typo in the main theorem.